# Exploiting Observation Bias to Improve Matrix Completion

### v2

@@ 4.1 Analysis
Proposition 4.3.
Let $\delta\in(0,1)$. Then, under the same assumption as in Proposition 4.2, the following event: $\forall(i,j)\in\mathcal{U}^{\star}\times\mathcal{V}^{\star}$,
@@
holds with probability at least $1-\delta$, where $c=\textup{poly}(d,B,L,m^{-1},\kappa^{-1})$ provided that $n\gtrsim d\log(e\,n/\delta)$, and $\varepsilon_{n}$ satisfies (89), and
@@
$\varepsilon_{n}$ $\lesssim\left(\frac{\kappa^{4}(m^{2}\wedge 1)}{dL(B\vee 1)}\wedge 1\right).$ (18)
The precise statement of Proposition 4.3 is given in Lemma D.2 and its proof in presented in Appendix D.
Prediction error.
Finally, we are ready to present the main performance guarantee satisfied by MNN. To simplify the exposition of our result, we will consider that $\rho_{n}=n^{-\beta}$ for some $\beta>0$.
Theorem 4.1.
Suppose we are under the same assumptions as in Proposition 4.2. Let $\delta\in(0,1)$. Suppose we choose $\varepsilon_{n}=2n^{-\frac{2-\beta}{d-1}}$. Then, the event: $\forall(i,j)\in[n]\times[n]$
$|\hat{x}_{ij}-x_{ij}|\leq K\,n^{-\frac{2-\beta}{2d}}\sqrt{\log\left(\frac{e\,n}{\delta}\right)},$ (19)
@@
$n\geq\left(\frac{CLBd^{3}}{m^{2}\kappa^{4}}\log\left(\frac{e\,n}{\delta}\right)\right)^{2d},$ (20)
where the constant $K=\textup{poly}(d,L,B,\sigma,m^{-1},\kappa^{-1})$, and $C>0$ is a universal constant.
The proof of Theorem 4.1 builds upon the previously stated propositions and its proof is deferred to Appendix F.
Observe that condition (20) in Theorem 4.1 can be verified [^1] This can be verified by using the principle branch of the Lambert function $W_{0}$, that satisfies $W_{0}(x)=y$ with $x=y^{y}$ for $x\geq 0$. Additionally, we have $W_{0}(x)\geq\log(x)-\log\log(x)+\frac{\log\log(x)}{\log(x)}$, for all $x\geq e$. (see Hoorfar and Hassani (2008)). if $d=O(\log(n)-\log(\log(n)))$ where $O(\cdot)$ may hide polynomial dependencies on $L,B,m^{-1}$ and $\kappa^{-1}$.
4.2 Discussion
Parametric rate in supervised learning.
Suppose we knew exactly the latent features $(\mathbf{u}_{i})_{i\in[n]}$, $(\mathbf{v}_{j})_{j\in[n]}$, then predicting outcomes of interest from noisy observations amounts to solving a supervised learning task where we recall that $f:\mathbb{R}^{2d}\mapsto\mathbb{R}$ is a bounded Lipschitz function that is a priori unknown (we recall from section 2 that we adopt a model of the form $y_{ij}=f(\mathbf{u}_{i},\mathbf{v}_{j})+\xi_{ij}$). In this classical supervised learning setting, we know, for example from Theorem 1 in Stone (1982), that the optimal achievable parametric prediction error rate is
$\max_{i,j\in[n]}|\hat{x}_{ij}-x_{ij}|=\widetilde{O}\left(N^{-\frac{1}{2+2d}}\right)$ (21)
where $N$ would correspond to the number of observations and $\widetilde{O}(\cdot)$ may hide logarithmic dependencies on $N$. Let us further add that the $k$-nearest neighbor is an example of supervised learning method that achieve such rate Chen and Shah (2018).
Comparison with the parametric rate.
@@ 4.2 Discussion
Lemma 4.1.
Let $N$ denote the expected number of observed noisy entries of $\mathbf{X}$. Suppose we are under the same assumptions as in Theorem 4.1, and with similar choices of $N_{n}$ and $\varepsilon_{n}$. Then for all $\delta\in(0,1)$, the event:
@@
holds with probability at least $1-\delta$, provided similar conditions as in Theorem 4.1 hold, and where $\widetilde{O}(\cdot)$ hides logarithmic dependencies on $\log(N)$ and polynomial dependencies in $d,L,B,\sigma,m^{-1},$ and $\kappa^{-1}$.
To spare the reader unnecessary details, we omit the proof of Lemma 4.1 as the result follows immediately form Theorem 4.1. To see that, note that under the proposed model, according to (2), the expected number of observation $N\approx n^{2-\beta}$. Rewriting this as $n=N^{-1/(2-\beta)}$ and plug-in it in the error rate expression (19) in Theorem 4.1 yields the desired result.
@@ 5.1 Synthetic Data
We performed several experiments with different numbers of datapoints $n$ and sparsity values $\rho_{n}$, all of which had data generated using the following steps. First, we generated rank $d=5$ latent factors $\mathbf{U}\in[-1,1]^{n\times d}$, $\mathbf{V}\in[-1,1]^{n\times d}$ by sampling $d$ entries i.i.d from a standard normal distribution for each of the $n$ users and items and then normalizing each row to have unit norm. We then generated the matrix $\mathbf{P}$ of probabilities as described in our observation model by setting $\mathbf{P}=\frac{\rho_{n}}{2}(\mathbf{U}\mathbf{V}^{\top}+\mathbf{1}\mathbf{1}^{\top})$. Next, we sample $\mathbf{A}\overset{\operatorname{iid}}{\sim}\mathsf{Ber}(\mathbf{P})$. We defined our outcome matrix $\mathbf{X}=\theta(\mathbf{U})\phi(\mathbf{V})^{\top}$ where
$\theta(\mathbf{U})=e^{\sqrt{d}\mathbf{U}}$ (23)
and
$\phi(\mathbf{V})=e^{\sqrt{d}\mathbf{V}}.$ (24)
We then generated the subgaussian noise terms $\xi_{ij}$ by sampling $n^{2}$ entries i.i.d from a standard normal distribution. Finally, the partially-observed noisy outcome matrix $\mathbf{Y}$ was generated by setting $y_{ij}=x_{ij}+\xi_{ij}$ if $a_{ij}=1$ and $y_{ij}$ to be a null value otherwise.
Algorithms compared.
@@
We compare the performance of the proposed algorithm against a standard matrix completion algorithm USVT. In the experiments, we utilize a practical implementation of the proposed MNN algorithm where (i) $\mathbf{A}$ is used in place of $\tilde{\mathbf{A}}$ in the distance estimation step and (ii) the matrix completion from the clustered outcome matrix is done using SVD with known rank. Using such a practical implementation of the algorithm acts a proof of concept – letting us determine the added value of the extra information gained from the two stage approach. We compare against a modified version of USVT where the rank of the matrix is known, allowing for more comparable performance against MNN which utilizes the matrix rank. We note that MNN is designed to handle biased data while the modified USVT algorithm is not.
Metrics.
@@
We use $R^{2}$ score, MSE, MAE, and max error to assess the similarity of the estimated full outcome matrix with the true outcome matrix. We directly compare the distribution of estimated entries to the distribution of true outcomes to assess the bias of the algorithms.
Experiment setup.
@@
We performed two different experiments with the described algorithms, using 16-fold cross-validation to tune the hyperparameters. In the first experiment, we tried 3 values of $n\in\{100,562,3162\}$. We set $\rho_{n}=n^{-1/4}$. We then ran both MNN and modified USVT with these values of $n$ and compared the resulting estimates. In the second experiment, we set $n=1000$ and tested 3 values of $\beta\in\{0,1/6,1/4\}$. We set $\rho_{n}=n^{-\beta}$. We then again ran both MNN and modified USVT and compared the bias of the resulting estimates. We concluded by running 10 experimental repeats of MNN and modified USVT using data generated with different random seeds for each of these $\rho_{n}=n^{-\beta}$ values. These final runs were used to calculate the $R^{2}$, MSE, MAE, and max error as a function of $n$ for each $\beta$.
Results.
@@
Before comparing the performance of MNN and modified USVT on the synthetic dataset, we examine the bias of the estimates for the full outcome matrix in both cases (experiments 1 and 2). As can be seen in Fig. 1,the distribution of estimates generated by MNN better approximates the true distribution of outcomes as the number of datapoints increases. Moreover, as $n$ increases, the shape of the distribution indicates that the results are less biased. However, the estimates generated by modified USVT do not display this same trend and do not appear to become less biased as the number of datapoints increases. This difference in bias can also be seen in Fig. 2 as $n$ is held constant at 1000. As $\rho_{n}$ decreases from 1 to $n^{-1/4}$, the bias of MNN barely changes but the bias of modified USVT increases significantly.
@@
Figure 1: Distributions for synthetic data generated with $n=100,562$, and $3162$ users and items (top to bottom).
@@
Figure 2: Predictions compared to true values for MNN (left) and modified USVT (right) with $\beta=0,\beta=1/6$, and $\beta=1/4$ (top to bottom).
@@
Table 1: $R^{2}$, MSE, MAE, and max error for matrix completion methods on synthetic datasets.
5.2 Real-World Data
Data.
We used a dataset obtained from Glance [^2] Glance (www.glance.com) is a smart lock screen that aims to personalize the screen. detailing the interaction of users with different pieces of content, considering a partially observed matrix where rows represent users, columns represent the different pieces digital content, and outcomes are the natural logarithm of the duration (in seconds) that the user interacted with the content. The dataset we considered had 1014696 observed measurements for 1305 users and 1471 content items. The sparsity of the dataset was thus $53\%$ and consistent with the observation model described in our setup.
Algorithms compared.
We compare the same practical implementation of MNN and a modified version of USVT described in the synthetic data experiments.
Metrics.
As in the synthetic data experiments, we used $R^{2}$ score, MSE, MAE, and max error to assess the similarity of our estimates with the true outcomes for the test sets. We also directly compare the distributions of estimated entries to true outcomes in our test set.
Experiment setup.
We split the data into $90/10$ train/test sets 10 times. We determined the best estimate of the rank of the true outcome matrix and the rank of the observation pattern using 9-fold cross-validation with MNN. We used 16-fold cross-validation to separately tune the other hyperparameters. Next, we ran both MNN and modified USVT on the training set in each experimental run. Using the tuned hyperparameters, we ran both algorithms for 10 experimental runs using different train/test splits to calculate performance metrics.

### v3

@@ 4.3 The completed matrix $\widehat{\mathbf{H}}$
Proposition 4.
Let $\delta\in(0,1)$. Under the setup of Proposition 3, the event: $\forall(i,j)\in\mathcal{U}^{\star}\times\mathcal{V}^{\star},$
@@
holds with probability at least $1-\delta$, provided that $n\gtrsim d\log(e\,n/\delta)$, and $\varepsilon_{n}$ satisfies (18), and
@@
$\varepsilon_{n}$ $\lesssim\frac{\kappa^{4}(m^{2}\wedge 1)}{dL(B\vee 1)}\wedge 1$ (20)
The proof of Proposition 4 is based on Proposition 2 and Proposition 3. We give its proof in Appendix C.
4.4 Main prediction error guarantee
Finally, we are ready to present the main performance guarantee satisfied by MNN. To simplify the exposition of our result, we will consider $\rho_{n}=n^{-\beta}$ for some $\beta>0$.
Theorem 1.
Let $\delta\in(0,1)$. Under the setup of Proposition 3, with choice the $\varepsilon_{n}=2n^{-(2-\beta)/(d-1)}$, the event:
$\max_{i,j\in[n]}|\hat{x}_{ij}-x_{ij}|\leq K\,n^{-\frac{2-\beta}{2d}}\sqrt{\log\left(\frac{e\,n}{\delta}\right)},$ (21)
@@ 4.4 Main prediction error guarantee
$n\geq\left(\frac{CLBd^{3}}{m^{2}\kappa^{4}}\log\left(\frac{e\,n}{\delta}\right)\right)^{2d}.$ (22)
Here, constant $K=\textup{poly}(d,L,B,\sigma,m^{-1},\kappa^{-1})$, and $C>0$ is a universal constant.
Observe that condition on $n$ in Theorem 1 can be verified [^4] We can verify by using the Lambert function $W_{0}$, which satisfies $W_{0}(x)=y$ with $x=y^{y}$ for $x\geq 0$. Additionally, we have $W_{0}(x)\geq\log(x)-\log\log(x)+\frac{\log\log(x)}{\log(x)}$, for all $x\geq e$. (see [20]). if $d=O(\log(n)-\log(\log(n)))$ where $O(\cdot)$ may hide polynomial dependencies on $L,B,m^{-1}$ and $\kappa^{-1}$. We shall now present the proof of Theorem 1 which is a direct consequence of Proposition 4 and Lemma 2.
Proof of Theorem 1.
Fix $i,j\in[n]$. We know, thanks to our clustering procedure, that there exist clusters $k\in\mathcal{U}^{\star}$, $\ell\in\mathcal{V}^{\star}$ such that $i\in\mathcal{U}_{k}$ and $j\in\mathcal{V}_{\ell}$. Thus, we may decompose the prediction error as follows
$x_{ij}-\hat{x}_{ij}=x_{ij}-h_{k\ell}+h_{k\ell}-\hat{h}_{k\ell}.$ (23)
Step 1. (Bounding the first term). First, we note that
$|h_{k\ell}-x_{ij}|$ $=|f(\mathbf{u}_{k},\mathbf{v}_{\ell})-f(\mathbf{u}_{i},\mathbf{v}_{j})|$ (24)
$\leq\sqrt{d}BL(\|\mathbf{u}_{k}-\mathbf{u}_{i}\|_{2}+\|\mathbf{v}_{\ell}-\mathbf{v}_{j}\|_{2}).$ (25)
We know, under our clustering procedure, that $\hat{d}_{ki}^{\mathbf{u}}\leq 6\varepsilon_{n}$ and $\hat{d}_{\ell j}^{\mathbf{v}}\leq 6\varepsilon_{n}$. Hence, using Lemma 2, we obtain that the event:
$|h_{k\ell}-x_{ij}|$ $\leq\sqrt{d}BL(\|\mathbf{u}_{k}-\mathbf{u}_{i}\|_{2}+\|\mathbf{v}_{\ell}-\mathbf{v}_{j}\|_{2})$ (26)
$\leq 14\sqrt{d}BL\varepsilon_{n}$ (27)
$n$ $\geq Cd\log\left(\frac{e\,n}{\delta}\right),$ (29)
where $c,C>0$ are universal constants, and recall that $\rho_{n}=n^{-\beta}$. Note that if we wish to obtain a non-trivial error rate, then condition (28) implies that we require $\beta<\frac{1}{4}.$
Step 2. (Bounding the second term). Second, we invoke Proposition 4, to obtain that the event:
$|\hat{h}_{k\ell}-h_{k\ell}|\leq c\left(\frac{d^{3}B^{2}}{(m^{2}\kappa^{4}\wedge m^{4}\kappa^{8})}\wedge 1\right)\left(LB\varepsilon_{n}+\frac{\sigma}{n}\left(\frac{2}{\varepsilon_{n}}\right)^{d-1}\sqrt{\frac{1}{\rho_{n}}\log\left(\frac{e\,n}{\delta}\right)}\right)$
$\left(\frac{\varepsilon_{n}}{2}\right)^{d-1}$ $\geq\frac{Cd}{m^{2}\kappa^{4}n}\left(\sqrt{\frac{1}{\rho_{n}}\log\left(\frac{e\,n}{\delta}\right)}+\log\left(\frac{e\,n}{\delta}\right)\right)$
$=\frac{Cd}{m^{2}\kappa^{4}n}\left(n^{\beta/2}\log^{1/2}\left(\frac{en}{\delta}\right)+\log\left(\frac{en}{\delta}\right)\right)$ (30)
where $C>0$ is some universal constant.
Clearly, choosing $\alpha=\frac{2-\beta}{2d},$ yields the minimal error rate
Next, we can also easily verify that condition (30) holds if
Finally, we easily verify that condition (31) holds if
In summary, we see that when
5 Comparison with lower bounds in supervised learning
Parametric rates in supervised learning.
Suppose we knew exactly the latent features $(\mathbf{u}_{i})_{i\in[n]}$, $(\mathbf{v}_{j})_{j\in[n]}$, then predicting outcomes of interest from noisy observations amounts to solving a supervised learning task where we recall that $f:\mathbb{R}^{2d}\mapsto\mathbb{R}$ is a bounded Lipschitz function that is a priori unknown (we recall from section 2 that we adopt a model of the form $y_{ij}=f(\mathbf{u}_{i},\mathbf{v}_{j})+\xi_{ij}$). In this classical supervised learning setting, we know, for example from Theorem 1 in [40], that the optimal achievable parametric prediction error rate is
$\max_{i,j\in[n]}|\hat{x}_{ij}-x_{ij}|=\widetilde{\Theta}\left(N^{-\frac{1}{2+2d}}\right)$ (40)
where $N$ would correspond to the number of observations and $\widetilde{O}(\cdot)$ may hide logarithmic dependencies on $N$. Let us further add that the $k$-nearest neighbor is an example of supervised learning method that achieves such rate, for example see [15].
Comparison with parametric rates in supervised learning.
@@ 5 Comparison with lower bounds in supervised learning
Lemma 1.
Consider the setup of Theorem 1. Let $N$ denote the expected number of observed noisy entries of $\mathbf{X}$. Then for all $\delta\in(0,1)$, the event
@@
holds with probability at least $1-\delta$, provided conditions of Theorem 1 hold. Here $\widetilde{O}(\cdot)$ hides logarithmic dependencies on $N$ and polynomial dependencies in $d,L,B,\sigma,m^{-1},$ and $\kappa^{-1}$.
We omit the proof of Lemma 1 as the result follows immediately from Theorem 1. To see that, note that under the proposed model, according to (2), the expected number of observation $N\approx n^{2-\beta}$. Rewriting this as $n\approx N^{-1/(2-\beta)}$ and plugging it in the error rate expression (21) in Theorem 1 yields the desired result.
@@
It is worth nothing that Theorem 1 assumes that $f$ factorizes as in (1) with number of factors being $d$. Indeed as $d$ scales it becomes a generic Lipschitz function but then the dependence on $d$ in Theorem 1 would start dominating. The assumption (1) is also the reason why $N^{-\frac{1}{2d}}\ll N^{-\frac{1}{2+2d}}$ is not a contradiction.
@@ 6 Experiments
6.1 Real-world data
In this section, we provide few experimental results on a real-world dataset to illustrate the performance gains promised by our theoretical findings.
Data. We used a dataset obtained from Glance [^5] Glance (www.glance.com) is a smart lock screen that aims to personalize the screen. detailing the interaction of users with different pieces of content, considering a partially observed matrix where rows represent users, columns represent the different pieces digital content, and outcomes are the natural logarithm of the duration (in seconds) that the user interacted with the content. The dataset we considered had 1014696 observed measurements for 1305 users and 1471 content items. The sparsity of the dataset was thus $53\%$ and consistent with the observation model described in our setup.
Algorithms compared. We utilize a practical implementation of the proposed MNN algorithm where (i) $\mathbf{A}$ is used in place of $\tilde{\mathbf{A}}$ in the distance estimation step and (ii) the matrix completion from the clustered outcome matrix is done using SVD with known rank. Using such a practical implementation of the algorithm acts a proof of concept – letting us determine the added value of the extra information gained from the two stage approach. The details of such implementation are provided in Section 6.4. We compare the performance of such implementation against a modified version of USVT, namely a version that has knowledge of the rank. We sought to further compare against SNN of [4], an algorithm achieving state-of-the-art results. However, due to prohibitive computational requirement, SNN is impractical for values of $n$ as large as in the Glance dataset. Instead, we compare SNN with MNN on smaller size synthetic data as discussed in Section 6.3.
Metrics. We use $R^{2}$ score, Mean Squared Error (MSE), Mean Absolute Error (MAE), and Maximum Across All Entries (max error) to assess performance of estimated outcome matrices. We directly compare the distribution of estimated entries to the distribution of true outcomes to assess the bias of the algorithms.
Experiment setup. We split the data into $90/10$ train/test sets at random and repeat the experiment 10 times. We determined the best estimate of the rank of the true outcome matrix and the rank of the observation pattern using 9-fold cross-validation with MNN. We used 16-fold cross-validation to separately tune the other hyperparameters. Next, we ran both MNN and modified USVT on the training set in each experimental run. Using the tuned hyperparameters, we ran both algorithms for 10 experimental runs using different train/test splits to calculate performance metrics.
Results. As we can see from Fig. 4, the estimates from modified USVT are extremely biased. The estimates from MNN, however, appear to be minimally biased and inline with ground truth. Moreover, from Fig. 3, we can see that the estimates made by modified USVT are very sensitive to outliers in the data, while the estimates from MNN are not. We also observe from Table 1 that the $R^{2}$, MSE, MAE, and max error of MNN are far better than that of modified USVT. Specifically the MSE of MNN is > 28x better compared to modified USVT. That is, MNN works significantly better on MNAR data.
Algorithm $R^{2}$ MSE MAE Max Error
USVT $-26.0\pm.435$ $32.4\pm.392$ $4.95\pm.032$ $11.6\pm.053$
Table 1: Comparison of performance of MNN and USVT on Glance data. As can be seen, MSE for MNN is >28x better.
Figure 3: Distributions of test set estimates from MNN and modified USVT for real-world data. The dashed line represents the mean of the true distribution. As can be seen MNN is reasonably accurate while USVT is large bias.
Figure 4: Predictions vs true outcomes for real-world data using MNN (left) and modified USVT (right). The dashed line represents perfect prediction. As can be seen MNN is reasonably accurate while USVT is far from the ground truth.
@@ 6.2 Synthetic Data
Data. We performed several experiments with different numbers of datapoints $n$ and sparsity values $\rho_{n}$, all of which had data generated using the following steps. First, we set $d=5$ and generated $d$-dimensional latent factors from the unit sphere for each of the $n$ users and items. Next, we generated the matrix $\mathbf{P}$ of probabilities and observation matrix $\mathbf{A}$ as described in our model. We defined the outcome matrix $\mathbf{X}=\theta(\mathbf{U})\phi(\mathbf{V})^{\top}$ where $\theta(\mathbf{U})=\exp({\sqrt{d}\mathbf{U}})$ and $\phi(\mathbf{V})=\exp({\sqrt{d}\mathbf{V}}).$ and $\exp(\cdot)$ is applied entry-wise. We then generated the subgaussian noise terms $\xi_{ij}$ by sampling $n^{2}$ entries i.i.d from a standard normal distribution. Finally, the partially-observed noisy outcome matrix $\mathbf{Y}$ was generated by setting $y_{ij}=x_{ij}+\xi_{ij}$ if $a_{ij}=1$ and $y_{ij}$ to be a null value otherwise.
@@
Algorithms compared. We compare the performance of the proposed algorithm against a standard matrix completion algorithm USVT. In the experiments, we utilize a practical implementation of the proposed MNN as explained in Section 6.4. We compare against a modified version of USVT where knowledge of the rank of the matrix is known, allowing for more comparable performance against MNN which utilizes the matrix rank. We note that MNN is designed to handle biased data while the modified USVT algorithm is not.
@@
Metrics. We use $R^{2}$ score, MSE, MAE, and max error to assess the similarity of the estimated full outcome matrix with the true outcome matrix. We directly compare the distribution of estimated entries to the distribution of true outcomes to assess the bias of the algorithms.
@@
Experiment setup. We performed two different experiments with the described algorithms, using 16-fold cross-validation to tune the hyperparameters. In the first experiment, we tried 3 values of $n\in\{100,562,3162\}$. We set $\rho_{n}=n^{-1/4}$. We then ran both MNN and modified USVT with these values of $n$ and compared the resulting estimates. In the second experiment, we set $n=1000$ and tested 3 values of $\beta\in\{0,1/6,1/4\}$. We set $\rho_{n}=n^{-\beta}$. We then again ran both MNN and modified USVT and compared the bias of the resulting estimates. We concluded by running 10 experimental repeats of MNN and modified USVT using data generated with different random seeds for each of these $\rho_{n}=n^{-\beta}$ values. These final runs were used to calculate the $R^{2}$, MSE, MAE, and max error as a function of $n$ for each $\beta$.
@@
Results. Before comparing the performance of MNN and modified USVT on the synthetic dataset, we examine the bias of the estimates for the full outcome matrix in both cases (experiments 1 and 2). As can be seen in Fig. 5,the distribution of estimates generated by MNN better approximates the true distribution of outcomes as the number of datapoints increases. Moreover, as $n$ increases, the shape of the distribution indicates that the results are less biased. However, the estimates generated by modified USVT do not display this same trend and do not appear to become less biased as the number of datapoints increases. This difference in bias can also be seen in Fig. 6 as $n$ is held constant at 1000. As $\rho_{n}$ decreases from 1 to $n^{-1/4}$, the bias of MNN barely changes but the bias of modified USVT increases significantly.
@@
Figure 5: Distributions for synthetic data generated with $n=100,562$, and $3162$ users and items (top to bottom). The dashed line represents the mean of the true distribution.
@@
Figure 6: Predictions compared to true values for MNN (left) and modified USVT (right) with $\beta=0,\beta=1/6$, and $\beta=1/4$ (top to bottom). The dashed line represents perfect prediction.
@@
Table 2: $R^{2}$, MSE, MAE, and max error for matrix completion methods on synthetic datasets (average $\pm$ standard deviation across 10 experimental repeats).
@@
6.3 Comparison with SNN
Experimental setup on the synthetic data.
We compared our method to SNN [4], a matrix completion algorithm for the MNAR setting. We only compared with SNN because it achieves state-of-the-art results. We compared the two algorithms on our synthetically generated data as described in section 6.2. The implementation of MNN is as described in Section 6.4. We considered $\beta=1/4$ and run the experiments with $n$ items and $n$ users where $n\in\{100,200,300,1000\}$.